OpenStudy (anonymous):

I GIVE GOLD MEDAL! (: The first four terms of a sequence is shown below. 9, 5, 1, -3 which of the following functions best defines this sequence?

OpenStudy (anonymous):

f(x)_1 = 9 f(x)_2 = f(x)_1 - 4 so your sequence will be 9, 5, 1, -3, -7, -11...

OpenStudy (anonymous):

so it'll be f(1) = 9, f( n + 1 ) = f(n) - 4, for n > 1 ? @Matt.Mawson

OpenStudy (anonymous):

correct!

OpenStudy (anonymous):

okay thank you!

OpenStudy (anonymous):

Thanks, that was fun (-: